OpenStudy (aliriodaza8):

Given f(x) and g(x) = f(x) + k, look at the graph below and determine the value of k. k = ______ Graph of f of x and g of x. f of x equals 1 over 3 x minus 2 and g of x equals 1 over 3 x plus 3.

OpenStudy (decarr432):

screen cap the graph and post it please

jhonyy9 (jhonyy9):

,,graph below" - where is ?

OpenStudy (dannyo19):

You could try setting this up as one equation with variables x and k

OpenStudy (dannyo19):

Then solve k in terms of x?

OpenStudy (scarlettfarra2000):

So we turn what they gave you into math-speak f(x)=1/3-2 g(x)=1/3+3 So we look between g(x) and f(x) k=g(x)-f(x) then subsitute f(x) = 1/3x-2 g(x) =1/3x+3 k=(1/3x+3) - (1/3x-2) then answer is 5

OpenStudy (mathmale):

I agree that the perpendicular distance between these two lines is 5. I'd strongly encourage you to sketch these two functions f(x) and g(x). doing so would immediately make this distance visual.

OpenStudy (mathmale):

alirio: If the original functions were presented symbolically, please stay with that representation, and don't change it to "Graph of f of x and g of x. f of x equals 1 over 3 x minus 2 and g of x equals 1 over 3 x plus 3." unless told to do so. Otherwise it's more work for you and for your helpers, and additional possibiities for errors arise.

OpenStudy (aliriodaza8):

thanks